A small body starts falling onto the Sun from a distance equal to the radius of the Earth orbit. The initial velocity of the body is equal to zero in the heliocentric reference frame. Making use of Kepler s laws, find how long the body will be falling.
1
Expert's answer
2020-10-19T13:25:24-0400

Solution

Here according to question semi major axis a=R/2

Hence "\tau" is the time of fall then

"(\frac{2\tau}{T}) ^2=(\frac{R/2}{R}) ^3"

By using kepler third law. 

So it's become

"\tau^2=\frac{T}{32}"

"\tau=\frac{365}{4\sqrt{2}}=64.5" days